OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. x2 - 64 = 0 A. -8 B. 16 C. -16 D. 2 E. 64 F. 8

OpenStudy (welshfella):

x^2 - 64 = 0 add 64 to both sides what do you get?

OpenStudy (welshfella):

-64 + 64 = ? 0 + 64 = ?

OpenStudy (anonymous):

64+64= 128 0+64= 64

OpenStudy (welshfella):

no -64 + 64 = 0 0 + 64 = 64 - thats correct

OpenStudy (welshfella):

so we have x^2 = 64 to find x we need to find the square roost of 64

OpenStudy (welshfella):

what number multiplied by itself gives 64?

OpenStudy (welshfella):

like 3 * 3 = 9 so 3 is the square root of 9 similarly 4 is the square root 16, as 4^2 = 16

OpenStudy (welshfella):

in fact there are 2 square roots because -3*-3 = 9 so the 2 square roots of 9 are 3 and -3 anf for 16 we have 4 and -4

OpenStudy (welshfella):

try multiplying some of the options by themselves and see if you can get 64 - if you do thats the answer